# Feasibility of Federated Learning from Client Databases with Different Brain Diseases and MRI Modalities

### v2

@@ Abstract
Segmentation models for brain lesions in MRI are commonly developed for a specific disease and trained on data with a predefined set of MRI modalities. Each such model cannot segment the disease using data with a different set of MRI modalities, nor can it segment any other type of disease. Moreover, this training paradigm does not allow a model to benefit from learning from heterogeneous databases that may contain scans and segmentation labels for different types of brain pathologies and diverse sets of MRI modalities. Additionally, the sensitivity of patient data often prevents centrally aggregating data, necessitating a decentralized approach. Is it feasible to use Federated Learning (FL) to train a single model on client databases that contain scans and labels of different brain pathologies and diverse sets of MRI modalities? We demonstrate promising results by combining appropriate, simple, and practical modifications to the model and training strategy: Designing a model with input channels that cover the whole set of modalities available across clients, training with random modality drop, and exploring the effects of feature normalization methods.  Evaluation on 7 brain MRI databases with 5 different diseases shows that such FL framework can train a single model that is shown to be very promising in segmenting all disease types seen during training. Importantly, it is able to segment these diseases in new databases that contain sets of modalities different from those in training clients.  These results demonstrate, for the first time, the feasibility and effectiveness of using Federated Learning to train a single 3D segmentation model on decentralised data with diverse brain diseases and MRI modalities, a necessary step towards leveraging heterogeneous real-world databases. Code will be made available at: https://github.com/FelixWag/FL-MultiDisease-MRI
Figure 1: FedUniBrain: Traditionally, brain lesion MRI segmentation models are trained for a specific disease, using MRI modalities determined by a study-specific protocol. Each resulting model is capable of only processing a predefined set of MRI input modalities, segmenting a single disease and is only trained on a limited number of samples. In this study, we introduce the FL framework FedUniBrain, which collaboratively trains a single model across multiple institutions capable of segmenting multiple diseases and is not bound to a predefined set of MRI input modalities. Our framework allows decentralized training without any data sharing, respecting the private nature of medical data.
@@ 1 Introduction
Brain MRI segmentation is crucial for disease analysis, diagnosis, and treatment. Multi-modal MRI scans provide complementary insights essential for detecting diseases. Each modality provides complementary information for tissues and anatomies via unique contrasts, which is crucial for diagnosing different diseases. For instance, the common image modality FLAIR suppresses cerebrospinal fluid signal and is therefore useful for detecting periventricular lesions like Multiple Sclerosis. Models that can jointly process all available MRI modalities for a patient are therefore paramount for brain MRI analysis. Traditionally, such multi-modal MRI models for segmentation of brain lesions are trained on data for a specific disease, acquired with a predefined protocol and set of MRI modalities [16, 17, 46, 44]. Models trained with this standard approach cannot segment other types of diseases, nor can they process patient scans if a different set of MRI modalities has been acquired compared to the specific set in the training data. While a recent study [41] demonstrated the feasibility of training a segmentation neural network across multiple brain MRI databases with different diseases and sets of modalities, that study used a centralised database as proof-of-concept. It did not address the practical challenge that, due to the sensitive nature of medical data and privacy concerns, data sharing between institutions is not straight-forward, which prevents the creation of an appropriately diverse centralised database. Federated Learning (FL) [27] allows collaboratively training a model without sharing or centrally aggregating data, protecting sensitive patient information. It is an open question whether it is possible to effectively train a model with FL using decentralised databases, each with different types of pathologies and sets of MRI modalities, which could then enable larger scale training of such models via multi-institutional collaborations.
@@
To address all these challenges, this study presents the FL framework FedUniBrain (Federated Universal Brain MRI Segmentation) for training a single 3D segmentation model on diverse brain lesion databases, each with a different disease and set of MRI input modalities, as shown in Fig. 1. This represents a challenging federated optimization setting since there is not only high heterogeneity within the MRI data from different scans across various databases but also across different segmentation tasks, as well as diverse MRI input modalities. While a few recent FL approaches have been developed to handle heterogeneous tasks and modalities per client, none of them is applied in the medical domain to address multiple MRI modalities per client for 3D segmentation. Central questions for this study are: Is federated training of such a model feasible? Will it be beneficial compared to training a disease-specific model? Will it result in a model that can segment new data with a different combination of modalities than the training data? Can a new client join in the training with a modality and/or pathology not seen during training?
@@
  • Feasibility: This study is the first to show feasibility of Federated training of a single segmentation model on decentralised databases with different brain diseases and sets of MRI modalities.
@@
  • Client with unseen modality and/or pathology joins the federation: We show that our framework is flexible and able to handle clients that join the federation with a previously unseen modality and/or pathology.
@@ 2 Related Work
Medical Image Segmentation on Heterogeneous Databases. Previous studies on medical image segmentation/lesion detection explored training with heterogeneous databases [42, 37, 28, 5]. However, these studies focus on centralised training, where data from different sources is combined into a single dataset. This overlooks practical real-world issues in the medical field, such as data sharing restrictions and the unique challenges associated with decentralized optimization. Additionally, they did not address the complexities of training with data with diverse sets of MRI modalities. A recent study [41] examined heterogeneous databases with diverse sets of MRI modalities in a centralized scenario. In contrast, our study focuses on the challenging setting of decentralized FL, which has not been explored yet.
Previous works on FL for brain MRI segmentation [20, 34, 29] focused on single-disease scenarios, where each client has access to the same set of MRI input modalities. These studies demonstrated that FL can effectively train a brain MRI segmentation without centralizing data. However, they did not address the challenges of diverse sets of MRI modalities and segmentation tasks per client.
@@
Federated Multi-Task Learning. Traditional Multi-Task learning trains a model on multiple tasks simultaneously to extract shared knowledge and generalize across multiple tasks [7, 11, 30]. Early research on Federated Multi-Task Learning mainly addresses data heterogenity across clients within a Multi-Task learning framework. Recent approaches [6, 24, 9] in the non-medical domain focus on both data heterogenity as well as task heterogenity per client (e.g. each client specializes on a different segmentation task). However, none of these works is applied to medical imaging and applicable to multi-MRI modalities inputs. In medical imaging, there exists only a few papers that tackle Multi-Task FL. [35] studied FL for learning from data with different labels in segmentation of CT. [14] tackles the problem of multi-task classification of mental disorders from MRI scans. However, both of theses works only consider one imaging modality as input. In summary, while there are works on FL with multi-modal MRI and FL for heterogeneous tasks; the combination of both, especially with different sets of modalities per client, has not been explored.
@@
Missing Modalities. Learning from databases with varying sets of MRI modalities relates to research of developing multi-modal models that perform well even when a modality is missing at inference time [12, 33, 36, 13, 47, 39]. Missing modalities in clinical practice are often caused by acquisition protocols, scanning costs or scanner availability. This line of work focuses on learning a model for a specific disease, from data with the same set of modalities during training. Furthermore, these methods are trained in centralised settings. FL for missing modalities is an emerging research direction. Most existing studies focus on text and image modalities [31, 4], non-imaging data [8], or require unimodal or complete multimodal data [43]. Notably, none of these approaches have been applied to different MRI modalities to segment multiple tasks. While addressing missing modalities is an important capability of our method, it is not the primary topic of this research. Instead, this study concentrates on the effectiveness and practicality of learning from diverse databases with different sets of input modalities and segmentation tasks.
@@
Comparison to Foundation Models. Our study has some similarities with the area of foundation models that are able to perform well and generalize to a variety of tasks, like SAM [18] and especially for medical imaging MedSAM [25]. However, there is a clear distinction to our work. Those models are not designed to handle 3D input data as well receive multi modality MRI input. Additionally, our method is not an interactive method that requires any user input prompts. Therefore, this line of research is not directly applicable to our problem setting.
@@ 3.1 Problem setting: FL with Multi-Modal MRI and Multiple Tasks
We consider $C$ clients, where each client $c$ ($1\!\leq\!c\!\leq\!C$) is associated with a multi-modal MRI dataset $\mathcal{D}_{c}=\{(\mathcal{X}_{ci},y_{ci})\}_{i=1}^{|\mathcal{D}_{c}|}$ with set of MRI-modalities $\mathcal{M}_{c}$, where $\mathcal{X}_{ci}=\{{x}_{ci}^{m}\}_{m=1}^{|\mathcal{M}_{c}|}$ denotes the multi-modal MRI image and $y_{ci}$ the corresponding segmentation mask of the $i$-th sample. Here, each $x_{ci}^{m}\in\mathbb{R}^{w\times h\times d}$, represents a 3D image, where $w,h$ and $d$ denote the width, height, and depth respectively. Furthermore, $\mathcal{M}=\bigcup_{i=1}^{C}M_{i}$ denotes the set that represents all unique modalities across all clients and $|\mathcal{M}|$ is the total number of unique modalities across all datasets. Each client solves a specific segmentation task $\mathcal{T}_{c}$. Our goal is to train, without data sharing, a single global model parameterized by $\theta$, that performs well on all tasks and clients. The global objective is to minimize the average of the client-local loss functions:
@@
Table 1: Comparison between FL settings: Previous FL settings are not designed to simultaneously handle different tasks per client, multiple MRI input modalities, missing modalities, and varying sets of modality combinations across client. FedUniBrain is designed to handle all these aspects simultaneously.
Setting Multi Task Multi MRI Mod. Missing Mod. Unseen Mod. Comb.
@@ 3.2 Unified Model for Multiple Segmentation Tasks
Client with unseen modality joins: In a realistic FL scenario, a new client may join a federation at a later training stage, for example, by contributing data to further improve a pre-trained model. If their data includes a new modality that was not in the originally available set $\mathcal{M}$, can we learn the new modality and data? To handle this scenario, we alter the model as follows. Assume a network pre-trained before the new client joins, and let the weights of its first convolutional layer be $\mathbf{W}_{1}\!\in\!\mathbb{R}^{o\times|\mathcal{M}|\times K_{h}\times K_{w}\times K_{d}}$, where $o$ is the number of output filters, $K_{h},K_{w},K_{d}$, the height, width, and depth of the kernel, respectively. To add the new modality, we extend dimensionality of $\mathbf{W_{1}}$ to get ${\mathbf{W^{\prime}_{1}}\in\mathbb{R}^{o\times(|\mathcal{M}|+1)\times K_{h}\times K_{w}\times K_{d}}}$. We then initialize the weights of the existing channels with the pre-trained weights: ${\mathbf{W^{\prime}_{1}}[:,i,:,:,:]=\mathbf{W_{1}}[:,i,:,:,:],\forall\,i\in[1,|\mathcal{M}|].}$ For the new modality, we randomly select one of the pre-existing modalities, sampling $k\in[1,|\mathcal{M}|]$, and use its pre-trained weights to initialize the weights associated with the newly added modality: ${\mathbf{W^{\prime}_{1}}[:,|\mathcal{M}|+1,:,:,:]=\mathbf{W_{1}}[:,k,:,:,:]}$. All network parameters, including $\mathbf{W^{\prime}_{1}}$, are then trained as normal with data from all clients.
@@
Algorithm: To summarise, our simple yet effective algorithm is shown in Fig. 2. Initially, the central server receives a list of available MRI modalities from each client’s dataset. Afterwards, the server initializes a model architecture that accepts the set of unique MRI modalities $\mathcal{M}$ as input, and the FL algorithm starts (Sec. 3.1). During training, each client’s input modalities are extended to the whole set of unique MRI modalities, with each missing modality filled with zeroes, and modality drop is applied.
@@ 4.2 Experimental Setup
All experiments use the Residual U-Net [45] as backbone, Adam optimizer, and $\alpha\!=\!0.8$ for our weighted loss function. We linearly scale the learning rate from 0 to $10^{-3}$ over 50 communication rounds, keep it constant until communication round 150, and then linearly decay it. We set $\tau\!=\!58$ (equivalent to 1 epoch for the largest database, ATLAS), $E\!=\!300$ (or 300 epochs for centralised training), and batch size of 8. The Dice score on the validation set is reported after the last epoch. For Federated Learning, we study the effect of feature normalisation methods, assessing BN when all its parameters are aggregated as per the standard FedAvg algorithm [27], and, separately, using FedBN [21] to keep client-specific BN parameters (cf. Sec. 3.2). We also assess methods that do not rely on client-specific normalization layers: GN [40], IN [38], and NF [48].
@@ 4.3 Results & Discussion
Experimental results in Tab. 2 prove that it is feasible to use Federated Learning to train a single model jointly over client databases with diverse diseases and sets of modalities and reach performance levels of database-specific models. When learning client-specific BN parameters, FedUniBrain matches or even surpasses single center training, highlighting the benefits of knowledge transfer between brain diseases. Training with modality drop and client-specific BN layers leads to the highest overall results and performs better than using averaged BN parameters, demonstrating the benefits of client specific BN layers for source clients.  Visual examples of training a single model per client compared to FedUniBrain are shown in Fig. 3(a). Besides its practicality over disease-specific models, the unified model may have improved generalisation abilities on unseen databases, which are assessed in follow-up experiments.
@@
Modality Drop for Robustness to Missing Modalities
Our desiderata is to train a model that can segment pathologies even in data with modality combinations different from those in the original training databases. This requires robustness against missing modalities. Results in Tab. 5 show that training with modality drop is effective for this purpose. When training with modality drop, the Dice score decreases by only 3.9% when evaluating with missing modalities compared to all modalities, versus 21.3% decrease without modality drop. Therefore, the following experiments focus on models trained with modality drop.
@@
Zero-Shot Generalization to Unseen Modality Combinations
We here assess the capabilities of generalizing to unseen databases with modality combinations not seen during training by evaluating on ISLES and Tumor2. Related results are shown in Tab. 4. Remarkably, despite not having seen any data from ISLES or Tumor2 during training, the best federated model’s average Dice score is only 0.9% lower than the Oracle models, which were trained on the specific target databases. FedUniBrain trained with FL across 5 clients reaches performance levels of centralized training using the 5 databases (MultiUnet, with IN as it performed best), demonstrating feasibility of decentralised training with such diverse brain lesion MRI databases.
Training with modality drop proves effective in dissociating specific modality combinations with certain pathologies. FedUniBrain with modality drop tends to show improved average performance over the two new databases with modality combinations not seen during training. Visual results are shown in Fig. 3(b). There is a large improvement for Tumor2 in all settings. Similarly, for ISLES, there is a general trend of improvement in most cases. In the few instances where the federated models performed slightly worse with modality drop, it may be due to stochasticity in federated optimization.
@@
Results reported in Tab. 2 indicate that learning client-specific parameters of BN yields better performance when evaluating on data from source clients, compared to other normalization methods. As shown in Tab. 4, however, when evaluating generalization on unseen databases (ISLES, Tumor2), BN with parameters tailored to source-clients underperforms compared to IN and NF. Consequently, a trade-off must be made when designing a model for training with this type of heterogeneous data: either prioritize performance on source clients that contributed training data or generalize to unseen databases.
@@
We evaluate the flexibility of our FedUniBrain framework in incorporating a new client that joins the federation at a later stage of training to further enhance the model with knowledge from new data. We consider three scenarios. First, a client with a modality combination not previously seen during training joins. Initially, the model is trained on ATLAS, BRATS, MSSEG, TBI, and WMH, then Tumor2 database joins (Fig. 4(a)). Second, a client with a modality not seen during earlier training joins, specifically the ISLES database that includes the unseen DWI modality (Fig. 4(a)). Lastly, a client with a modality and pathology not seen during training joins. We test this scenario in two cases: excluding TBI from initial training, the only database with SWI modality and Traumatic Brain Injuries (Fig. 4(b)), and excluding MSSEG, which is the only database with PD modality and Multiple Sclerosis (Fig. 4(c)).
The FedUniBrain model is trained for 300 communication rounds before a new client joins, followed by 200 additional rounds (optimization as per 4.2). Results in Fig. 4 show that the Dice scores for the 5 initial training databases remain stable after a new client joins and even slightly improve when a client with a previously unseen pathology joins (Fig. 4(b)4(c)). Moreover, FedUniBrain learns to segment the new modalities and pathologies of new clients, outperforming training only on the specific client (single-center), which indicates positive knowledge transfer. Even in the challenging case when a client with a previously unseen modality joins, our framework handles it without training re-initialization or complex modifications, enabling seamless continuation of training. Experiments on zero-shot generalization (Tab. 4) showed strong performance on unseen databases with pathologies seen during training. However, as shown by the blue bars in Fig. 4(b) and 4(c), generalization struggles with pathologies not seen during training. The FedUniBrain framework can easily integrate new clients, even with previously unseen pathologies and modalities, and learn to effectively segment the new type of data.

### v3

@@ Abstract
Segmentation models for brain lesions in MRI are typically developed for a specific disease and trained on data with a predefined set of MRI modalities. Such models cannot segment the disease using data with a different set of MRI modalities, nor can they segment other types of diseases. Moreover, this training paradigm prevents a model from using the advantages of learning from heterogeneous databases that may contain scans and segmentation labels for different brain pathologies and diverse sets of MRI modalities. Additionally, the confidentiality of patient data often prevents central data aggregation, necessitating a decentralized approach. Is it feasible to use Federated Learning (FL) to train a single model on client databases that contain scans and labels of different brain pathologies and diverse sets of MRI modalities? We demonstrate promising results by combining appropriate, simple, and practical modifications to the model and training strategy: Designing a model with input channels that cover the whole set of modalities available across clients, training with random modality drop, and exploring the effects of feature normalization methods.  Evaluation on 7 brain MRI databases with 5 different diseases shows that this FL framework can train a single model achieving very promising results in segmenting all disease types seen during training. Importantly, it can segment these diseases in new databases that contain sets of modalities different from those in training clients.  These results demonstrate, for the first time, the feasibility and effectiveness of using FL to train a single 3D segmentation model on decentralised data with diverse brain diseases and MRI modalities, a necessary step towards leveraging heterogeneous real-world databases. Code: https://github.com/FelixWag/FedUniBrain
footnotetext: Published as a conference paper at WACV 2025.
Figure 1: FedUniBrain (yellow box) enables the collaborative training of a single brain MRI segmentation model capable of segmenting multiple brain pathologies across multiple institutions without data sharing. FedUniBrain can train a model on data from institutions with varying sets of input MRI modalities and different brain diseases. In contrast, standard brain MRI segmentation models (grey box), can only process MRI scans with a specific set of MRI input modalities and are disease-specific.
@@ 1 Introduction
Federated Learning (FL) [31] is a promising approach for training machine learning models in healthcare, where patient privacy and data-sharing restrictions are critical. In medical imaging, multi-institutional collaborations are essential for training models that perform well across diverse patient populations [40, 33]. However, privacy regulations prevent institutions from sharing confidential patient data, preventing the creation of an appropriately diverse centralised database. FL enables collaboratively model training across institutions without data sharing, allowing the use of larger, decentralized data. However, FL faces the challenges of data heterogeneity from different institutions, including variations in scanners, and disease specializations, complicating the decentralized optimization [24, 19]. This study focuses on multi-modal MRI segmentation, introducing additional heterogeneity in input modalities. Each MRI modality (e.g. FLAIR, T1, T2) provides complementary information about tissues and anatomies, which is important for diagnosing different diseases. For example, FLAIR is useful for detecting periventricular lesions in Multiple Sclerosis, T1c highlights enhancing tumour, while SWI highlights bleeding. Depending on equipment, disease specialization, and clinical protocols, institutions collect scans with different input modalities and brain pathologies. Therefore, training a unified MRI segmentation neural network across institutions with different sets of MRI modalities and diseases via FL introduces the challenge of handling varying sets of input modalities per institution, while being effective in segmenting multiple diseases. An additional challenge is enabling new clients joining the federation with previously unseen modality combinations and pathologies. Traditionally, multi-modal MRI models for brain lesion segmentation are trained on specific disease data, acquired with a predefined protocol and set of MRI modalities [18, 20, 51]. These models cannot segment other types of diseases, nor can they process scans with a different set of MRI modalities compared to the specific set in the training data. It is an open question whether it is possible to effectively train a model with FL using decentralised databases, each with different types of pathologies and sets of MRI modalities.
@@
To address these challenges, this study presents, FedUniBrain (Federated Universal Brain MRI Segmentation) a FL framework for training a single 3D segmentation model on diverse brain lesion databases, each with a different disease and set of MRI modalities (Fig. 1). This challenging setting involves three types of heterogeneities within the MRI data: difference in scans, different segmentation tasks (brain pathologies), and diverse sets of MRI input modalities across clients. While many FL works have addressed heterogeneous medical imaging data, this is the first one to address all three types of heterogeneities combined, whereas previous works mainly focused on one type and assumed the same set of modalities per client, making it unclear how to even train a single model with varying input modalities per clients. Central questions for this study are: Is federated training of such a model feasible? Will it be beneficial compared to training disease-specific models? Will it result in a model that can segment new data with a different combination of modalities than the training data? Can a new client join the training with a modality and/or pathology not seen during training?
@@
  • Feasibility: This study is the first to show feasibility of Federated training of a single segmentation model across decentralised databases, with varying sets of input modalities and segmentation tasks
@@
  • Client with unseen modality and/or pathology joins the federation: We show that our framework is flexible and can handle clients that join the federation with a previously unseen modality and/or pathology.
@@ 2 Related Work
Federated Learning on Heterogeneous Databases. Many previous studies have explored and demonstrated the potential of FL for heterogeneous medical imaging data [42, 23, 38, 32, 49, 11, 46, 17, 48, 13]. However, these works typically address heterogeneity with respect to scanner differences or label distribution, commonly using one imaging modality or the same set of modalities at each client. Previous works on FL specifically for brain MRI segmentation [23, 38, 32] focused on single-disease scenarios, where all clients have access to the same set of MRI input modalities. These studies demonstrated that FL can effectively train a multi-modal brain MRI segmentation model without centralizing data. However, none of these methods are applicable when clients have diverse sets of input modalities. Only recently, a study [45] demonstrated for the first time the feasibility of training a segmentation neural network across multiple brain MRI databases with different diseases and sets of modalities, but in a centralised settings, it used a centralised database as a proof-of-concept. It did not address the practical challenge of medical data sharing restriction. In contrast, our study focuses on the challenging setting of decentralized FL, which has not been explored yet.
@@
Federated Multi-Task Learning. Centralised Multi-Task learning trains a model on multiple tasks to extract shared knowledge and generalize across tasks [6, 10, 34]. Recent Federated Multi-Task Learning [5, 27, 8] focuses on both data and task heterogeneity in a federated setting (e.g. each client specializes on a different segmentation task), but none of these works is applied to medical imaging or applicable to multi-MRI modalities inputs. In medical imaging, few papers explore Multi-Task FL. [39] studies FL for learning from data with different segmentation labels in CT. [15] tackles the problem of multi-task classification of mental disorders from MRI scans, but both of theses works only use one imaging modality as input. In summary, while FL with multi-modal MRI and FL for heterogeneous tasks have been studied; the combination of both, especially with different sets of modalities per client, has not been explored.
@@
Missing Modalities. Learning from databases with varying sets of MRI modalities relates to research on missing modalities during inference [12, 37, 14, 52, 43], often caused by clinical factors like acquisition protocols, costs or scanner availability. Most of these methods focus on disease-specific models with the same modalities used during training and in a centralized setting. FL with missing modalities is a recent research area but mainly focusing on text and image modalities [35, 4], non-imaging data [7], or require unimodal or complete multimodal data [47]. Notably, none of these approaches is applicable to varying MRI modalities to segment multiple tasks. While addressing missing modalities is an important capability of our method, it is not the primary topic of this research.
@@
Comparison to Foundation Models. Our study shares similarities with foundation models like SAM [21] and MedSAM [28] for medical imaging, which perform and generalize well to a variety of tasks. However, these models are not designed for 3D input data or multi-modality MRI input. Additionally, unlike our method, these models are interactive and require user input prompts. Therefore, this research is not directly applicable to our problem.
@@ 3.1 Problem setting: FL with Multi-Modal MRI and Multiple Tasks
We consider $C$ clients, where each client $c$ ($1\!\leq\!c\!\leq\!C$) has a multi-modal MRI dataset $\mathcal{D}_{c}=\{(\mathcal{X}_{ci},y_{ci})\}_{i=1}^{|\mathcal{D}_{c}|}$ with the set of MRI-modalities $\mathcal{M}_{c}$, where $\mathcal{X}_{ci}=\{{x}_{ci}^{m}\}_{m=1}^{|\mathcal{M}_{c}|}$ denotes the multi-modal MRI image and $y_{ci}$ the corresponding segmentation mask for sample $i$. Here, each $x_{ci}^{m}\in\mathbb{R}^{w\times h\times d}$, represents a 3D image, where $w,h$ and $d$ denote the width, height, and depth respectively. Furthermore, $\mathcal{M}=\bigcup_{i=1}^{C}M_{i}$ denotes the set that represents all unique modalities across all clients and $|\mathcal{M}|$ is the total number of unique modalities across all datasets. Each client solves a specific segmentation task $\mathcal{T}_{c}$. Our goal is to train, without data sharing, a single global model parameterized by $\theta$, that performs well on all tasks and clients. The global objective is to minimize the average of the client-local loss functions:
@@
Table 1: Previous FL settings are not designed to simultaneously handle different tasks, multiple MRI input modalities, missing modalities, and varying sets of modality combinations across clients. FedUniBrain handles all these aspects simultaneously.
Setting Multi Task Multi MRI Mod. Missing Mod. Varying Mod. Comb.
@@ 3.2 Unified Model for Multiple Segmentation Tasks
Algorithm: An overview of our simple yet effective algorithm is shown in Fig. 2. Initially, the central server receives a list of available MRI modalities from each client. Afterwards, the server initializes a model architecture that accepts the set of unique MRI modalities $\mathcal{M}$ as input, and the FL algorithm starts (Sec. 3.1). During training, each client’s input modalities are extended to the whole set of unique MRI modalities, with each missing modality filled with zeroes, and modality drop is applied. Supplementary Alg. 1 contains the full algorithm for further details.
@@
Client with unseen modality joins: In a realistic FL scenario, a new client may join a federation at a later training stage, by contributing data to further improve a pre-trained model. If their data includes a new modality that was not in the originally available set $\mathcal{M}$, can we learn the new modality and data? To handle this scenario, we make the following model modification to process the new modality concatenated along with already known modalities as multi-channel input images. Assume a pre-trained network before the new client joins, and the first convolutional layer weights $\mathbf{W}_{1}\!\in\!\mathbb{R}^{o\times|\mathcal{M}|\times K_{h}\times K_{w}\times K_{d}}$, where $o$ is the number of output filters, $K_{h},K_{w},K_{d}$, the height, width, and depth of the kernel, respectively. To add the new modality, we extend dimensionality of $\mathbf{W_{1}}$ to get ${\mathbf{W^{\prime}_{1}}\in\mathbb{R}^{o\times(|\mathcal{M}|+1)\times K_{h}\times K_{w}\times K_{d}}}$. We then initialize the weights of the existing channels with the pre-trained weights: ${\mathbf{W^{\prime}_{1}}[:,i,:,:,:]=\mathbf{W_{1}}[:,i,:,:,:],\forall\,i\in[1,|\mathcal{M}|].}$ For the new modality, we randomly select one of the pre-existing modalities, sampling $k\in[1,|\mathcal{M}|]$, and use its pre-trained weights to initialize the weights associated with the newly added modality: ${\mathbf{W^{\prime}_{1}}[:,|\mathcal{M}|+1,:,:,:]=\mathbf{W_{1}}[:,k,:,:,:]}$. All network parameters, including $\mathbf{W^{\prime}_{1}}$, are then trained as normal with data from all clients.
@@ 4.2 Experimental Setup
All experiments use the Residual U-Net [50] as backbone, Adam optimizer, and $\alpha\!=\!0.8$ for our weighted loss function. We linearly scale the learning rate from 0 to $10^{-3}$ over 50 communication rounds, keep it constant until communication round 150, and then linearly decay it. We set $\tau\!=\!58$ (equivalent to 1 epoch for the largest database, ATLAS), $E\!=\!300$ (or 300 epochs for centralised training), and batch size of 8. The Dice score on the validation set is reported after the last epoch. For FL, we study the effect of feature normalisation methods, assessing BN when all its parameters are aggregated as per the standard FedAvg algorithm [31], and, separately, using FedBN [24] to keep client-specific BN parameters (cf. Sec. 3.2). We also assess methods that do not rely on client-specific normalization layers: GN [44], IN [41], and NF [53]. To strengthen our findings, we tested the statistical significance of the main results for superiority or non-inferiority, as appropriate. A detailed analysis can be found in the Supplementary (Sec. D).
@@ 4.3 Results & Discussion
Experimental results in Tab. 2 prove that it is feasible to use Federated Learning to train a single model jointly over client databases with diverse diseases and sets of modalities and reach performance levels of database-specific models. When learning client-specific BN parameters, FedUniBrain matches or even surpasses single center training, with statistical superiority for ATLAS (p=0.03) and WMH (p$<$0.001) and non-inferiority for BRATS (p$<$0.001), MSSEG (p=0.011), and TBI (p$<$0.001), highlighting the benefits of knowledge transfer between brain diseases. Training with modality drop and client-specific BN layers leads to the highest overall results and performs better than using averaged BN parameters, demonstrating the benefits of client specific BN layers for source clients.  Visual examples of training a single model per client compared to FedUniBrain are shown in Fig. 3(a). Besides its practicality over disease-specific models, the unified model may have improved generalisation abilities on unseen databases, which are assessed in follow-up experiments.
Modality Drop for Robustness to Missing Modalities
Our goal is to train a model that can segment pathologies even when modality combinations differ from those in the original training databases. This requires robustness against missing modalities. Results in Tab. 5 show that training with modality drop is effective for this purpose. When training with modality drop, the Dice score decreases by only 3.9% when evaluating with missing modalities compared to all modalities, versus 21.3% decrease without modality drop. Therefore, the following experiments focus on models trained with modality drop.
@@
Zero-Shot Generalization to Unseen Modality Combinations
We assess the generalization capabilities to unseen databases with modality combinations not seen during training by evaluating on ISLES and Tumor2. Related results are shown in Tab. 4. Remarkably, despite not having seen any data from ISLES or Tumor2 during training, the best federated model’s average Dice score is only 0.9% lower than the Oracle models, which were trained on the specific target databases. FedUniBrain trained with FL and IN across 5 clients reaches performance levels of centralized training using the 5 databases (MultiUnet, with IN as it performed best), which is confirmed by our statistical non-inferiority tests on ISLES (p=0.003) and Tumor2 (p$<$0.001). This demonstrates feasibility of decentralised training with such diverse brain lesion MRI databases.
Training with modality drop effectively disassociates specific modality combinations with certain pathologies. FedUniBrain with modality drop tends to show improved average performance over the two new databases with modality combinations not seen during training with visual results shown in Fig. 3(b). There is a large improvement for Tumor2 in all settings. Similarly, for ISLES, there is a general trend of improvement. Stochasticity in federated optimization may explain the few cases where modality drop made federated models perform slightly worse.
@@
Results in Tab. 2 indicate that learning client-specific BN parameters improves performance on source clients, compared to other normalization methods. As shown in Tab. 4, however, when evaluating generalization on unseen databases (ISLES, Tumor2), BN with parameters tailored to source-clients underperforms compared to IN and NF. Consequently, a trade-off must be made when designing a model for training with this type of heterogeneous data: either prioritize performance on source clients that contributed training data or generalize to unseen databases.
@@
We evaluate the flexibility of our FedUniBrain framework in incorporating a new client that joins the federation at a later stage of training to further enhance the model with knowledge from new data. We consider three scenarios. First, a client with a modality combination not seen during training joins. Initially, the model is trained on ATLAS, BRATS, MSSEG, TBI, and WMH, then Tumor2 database joins (Fig. 4(a)). Second, a client with a modality not seen during earlier training joins, specifically the ISLES database that includes the unseen DWI modality (Fig. 4(a)). Lastly, a client with a modality and pathology not seen during training joins. We test this scenario in two cases: excluding TBI from initial training, the only database with SWI modality and Traumatic Brain Injuries (Fig. 4(b)), and excluding MSSEG, which is the only database with PD modality and Multiple Sclerosis (Fig. 4(c)).
The FedUniBrain model is trained for 300 communication rounds before a new client joins, followed by 200 additional rounds (optimization as per 4.2). Results in Fig. 4 show that the Dice scores for the 5 initial training databases remain stable after a new client joins and even slightly improve when a client with a previously unseen pathology joins (Fig. 4(b)4(c)). Moreover, FedUniBrain learns to segment the new pathologies and modalities of new clients, outperforming training only on the specific client (single-center), as confirmed by our statistical superiority tests for pathologies seen during training ISLES (p=0.042) and Tumor2 (p=0.047), which indicates positive knowledge transfer. For the challenging setting of completely unseen modalities and pathologies of TBI (p$<$0.001) and MSSEG (p=0.006), non-inferiority tests show that FedUniBrain matches single-center performance. This demonstrates that our framework handles new clients seamlessly, without re-initialization or complex modifications, and not overfitting to previously learnt diseases. Experiments on zero-shot generalization (Tab. 4) show strong performance on unseen databases with pathologies seen during training. However, as shown by the blue bars in Fig. 4(b) and 4(c), generalization struggles with pathologies not seen during training. The FedUniBrain framework can easily integrate new clients, even with previously unseen pathologies and modalities, and learn to effectively segment the new type of data.